Of 30 students in a class, 15 passed. What percentage of the students passed?
Answer by stanbon(75887)

15/30 = 1/2 = = 0.50 = 50%
---------
Or you could do it this way:
15/30 = x/100
30x=1500
x= 50
Therefore 15/30 = 50/100 = 50%
-------------
Cheers,
Stan H.